# Unified approach to the resources of tensor network and stabilizer simulations

### v2

@@ Abstract
We introduce a general resource indicator, called the bra-ket entanglement, which can be used to bound the resource dependence of classical simulations in the tensor network framework and in the stabilizer formalism. For the tensor network framework, our bounds indicate that bra-ket entanglement governs the interplay between two physical resources, the coherence and the magic. As bra-ket entanglement increases, the dominant resource that governs the complexity of the tensor network framework, quantified by entanglement, shifts from coherence to magic. For the stabilizer formalism approach, we find that magic is always the dominant resource regardless of bra-ket entanglement. This conclusion is obtained by developing an operator stabilizer formalism, which extends the standard stabilizer formalism for pure states and has additional advantages in simulating certain quantum circuits. Therefore, our results indicate that as bra-ket entanglement increases, the resource governing the complexity of the two approaches go from different to the same.
Introduction.— Classical simulations of quantum systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] are a central tool for both science and technology. They set a benchmark for quantum computational advantages, and offer a powerful lens for exploring the boundary between classical and quantum computing [15, 16, 17, 18, 19, 20, 21]. Two major approaches to classical simulations are the tensor network [3, 4, 5] and the stabilizer frameworks [1, 2]. These two approaches have generally different domains of application, due to different physical resources determining the complexity of classical simulations. In the tensor network framework, the simulation complexity is governed by the bond dimension, which in the pure state case is connected to entanglement of the states to be simulated [22, 23]. In the stabilizer framework, instead, the simulation complexity is connected to another physical resource, known as “magic,” and mathematically associated to generating linear combinations of Pauli operators [2, 24, 25].
The different physical resources responsible for the complexity of simulations in the tensor network and stabilizer frameworks imply that states that are easy to simulate in one approach may be difficult to simulate in the other. For example, stabilizer states can be efficiently simulated in the stabilizer framework, but can have maximal bipartite entanglement [26], thereby becoming challenging for tensor networks. In contrast, low-bond matrix product states can be efficiently simulated with tensor networks, but can have large magic [27, 28], which makes them hard to simulate with the stabilizer framework. Finally, in the intermediate regime where both entanglement and magic are not too high, a combination of the tensor network and the stabilizer approach has been found to be beneficial [29, 30, 31, 32, 33].
Characterizing the interplay between entanglement and magic is therefore important, because it determines the choice of the most efficient classical simulation method for a given quantum experiment. Recently, relations between magic and operator entanglement [34] were observed for observables generated by unitary evolution from products of Pauli observables [35]. However, the relation between entanglement and magic is not clear in the general case, and even determining these quantities directly from the description of states and observables may not be straightforward.
In this paper, we introduce a quantity that relates entanglement and magic, providing useful bounds on the resources that governs the complexity for classical simulations of quantum states and observables in the tensor network framework and the stabilizer formalism. This quantity, which we call bra-ket entanglement, is defined for general operators $O$, and applies to the problem of simulating an operator of the form $UOU^{\dagger}$, where $U$ is a unitary operator. Here, $O$ can represent either a quantum state or a quantum observable, and $U$ represents the quantum evolution in the Schrödinger or in the Heisenberg picture, respectively. In the case of observables, this setting is important for the study of quantum chaos and information scrambling [36, 37].
We show that the amount of quantum entanglement in the operator $UOU^{\dagger}$ is determined by its magic and its coherence in the computational basis [38]. These two resources are in a tradeoff relation controlled by the the bra-ket entanglement of the operator $O$: as the bra-ket entanglement increases, the leading resource transitions from coherence to magic. This finding provides an immediate insight into the quantum resources determining the complexity of classical simulations in the tensor network approach, which we further corroborate by numerical experiments.
Our results indicate that magic is the leading resource that governs entanglement in the case of states and observables with high bra-ket entanglement, and provide an explanation for the correlations between entanglement and magic observed in Ref.[35]. In this regime, the tensor network and stabilizer approaches are comparable, as their complexity is affected by the same quantum features. Instead, in the low bra-ket entanglement regime, the hardness of classical simulation in the tensor network and the stabilizer formalism appears to have a different origin. In the stabilizer case, the results for $O$ beyond pure states are obtained by extending the standard stabilizer formalism to operators, developing an operator stabilizer formalism, and showing that magic is the key resource determining the simulation complexity regardless of bra-ket entanglement. We also show that the operator stabilizer formalism is much more powerful in simulating certain quantum circuits over the standard stabilizer formalism.
@@
Spatial vs bra-ket entanglement.— Let $O$ be an operator acting on the Hilbert space of $n$ qubits, hereafter denoted by ${\cal H}_{n}:=\mathbb{C}^{2^{n}}$, and let $|O\rangle:=\sum_{i,j}\,\langle i|O|j\rangle/\sqrt{\mathrm{Tr}[O^{\dagger}O]}\,|i\rangle|j\rangle$ be the corresponding normalized vector in the tensor product Hilbert space ${\cal H}_{n}\otimes{\cal H}_{n}$. In the following, we will take $O$ to represent either the density matrix of $n$ qubits, or an observable of the $n$-qubit system.
@@
where $H_{\alpha}(\{p_{i}\}):=\log\left(\sum_{i}p_{i}^{\alpha}\right)/(1-\alpha)$ is the $\alpha$-order Rényi entropy [39]. We call $H_{SE,\alpha}\left(O\right)$ the space entanglement (SE) of the operator $O$. When $O$ is the density matrix of a pure quantum state, its SE is exactly twice the entanglement entropy of the state [22]. When the operator $O$ is the density matrix of a mixed state, the spatial entanglement is generally not a measure of mixed state entanglement (for example, it can be nonzero even for some separable states) [40, 41], but it is important because it determines the bond dimension, which in turn determines the complexity of classical simulations in the tensor network framework [3]. In the literature, $H_{SE,\alpha}\left(O\right)$ is sometimes known the operator entanglement of operator $O$ [34].
@@
where $\{\gamma_{i}\}$ are the Schmidt coefficients in the decomposition $|O\rangle=\sum_{i}\gamma_{i}|O_{r,i}\rangle\otimes|O_{c,i}\rangle$, where $|O_{r,i}\rangle\}_{i}$ ($\{|O_{c,i}\rangle\}_{i}$) are orthogonal vectors acting on the Hilbert spaces associated to the row (column) indices (See SM B.2). For example, when $O$ is a rank-one operator like a pure state density matrix, $|O\rangle$ is a product state with zero BKE. In contrast, when $O$ is an $n$-qubit Pauli operator, $|O\rangle$ is proportional to a Bell state and has maximal BKE. In later sections, we will see how BKE relates the SE with two other quantum resources, namely magic and coherence.
Magic and coherence in the vectorization picture. Magic quantifies the deviation from stabilizer states, and the cost of simulating non-stabilizer states in the stabilizer framework [1, 2]. Here, we define a quantifier of magic in the vectorization picture. Our approach is to expand the vector $|O\rangle$ in the Bell basis with respect to the bipartition $(ab|cd)$ between bras and kets in the original operator $O$, and to define the Bell basis coherence (BBC) as the $\alpha$-order Rényi entropy
@@
Bra-ket bipartition diagnoses SE.—We define three free circuit families. The CBC-free circuit family $\mathcal{F}_{CBC}$ and BBC-free circuit family $\mathcal{F}_{BBC}$ are defined as sets of unitaries $U$ satisfying $H_{CBC,\alpha}(O)=H_{CBC,\alpha}(UOU^{\dagger})$ and $H_{BBC,\alpha}(O)=H_{BBC,\alpha}(UOU^{\dagger})$, respectively, for arbitrary $|O\rangle$. $\mathcal{F}_{CBC}$ is generated by diagonal phase and permutation gates on the computational basis, such as $\{T,S,CX\}$ or $\{CCX,S,CX\}$, while $\mathcal{F}_{BBC}$ consists of Clifford circuits generated by $\{H,S,CX\}$. The zero-coherence circuit family $\mathcal{F}_{zero}$ is the intersection of $\mathcal{F}_{CBC}$ and $\mathcal{F}_{BBC}$ with circuits constructed from $\{S,CX\}$ as examples. We discuss similarities and differences between $\mathcal{F}_{CBC}$ and $\mathcal{F}_{BBC}$ in SM E.1.
@@
We address this question in two steps. The first step is to relate SE with CBC and BBC.
@@
Detailed proofs can be found in SM D.1. Note that CBC and BBC are respectively sensitive to Hadamard and $T$ gates, which are responsible for coherence and magic generation, respectively. For a circuit $U$ with $N_{H}$ Hadamards and $N_{T}$ $T$ gates, we have (see SM C.1 and C.2):
@@
Therefore, in the absence of additional CBC and BBC resources ($H$ and $T$), $G_{zero,\alpha}(O)$ is governed by the existing CBC and BBC in $|O\rangle$. To generate higher entanglement, additional CBC and BBC resources are needed, especially the smaller one. Theorem 1 explains why the stabilizer states can achieve maximal entanglement. Starting from $O=|0\rangle\langle 0|^{\otimes n}=\prod_{i=1}^{n}(I_{n}+Z_{i})/2$, $H_{CBC,\alpha}(O)=0$ severely limits $G_{zero,\alpha}(O)$, therefore, $H$ in $\mathcal{F}_{BBC}$ are necessary to increase RHS of Theorem 1. On the other hand, we have $H_{BBC,\alpha}(O)=n$ already reaches the largest achievable SE, meaning no need to introduce additional magic resources. Therefore, we have $0=G_{zero,\alpha}\left(|0\rangle^{\otimes n}|0\rangle^{\otimes n}\right)=G_{CBC,\alpha}\left(|0\rangle^{\otimes n}|0\rangle^{\otimes n}\right)$ while $G_{BBC,\alpha}\left(|0\rangle^{\otimes n}|0\rangle^{\otimes n}\right)=n$. Similar reasoning applies to $O=P$, with opposite behaviors. In fact, Theorem 1 can be generalized to coherence defined on arbitrary product bases, whose proof is given in Lemma 6 of SM D.1.
The second step is to understand what determines CBC and BBC. From above, we observed that for $O=|0\rangle\langle 0|^{\otimes n}$ with minimal BKE, CBC is required to generate higher entanglement, while for $O=P$, with maximal BKE, BBC is necessary to generate higher entanglement. Therefore, the BKE of $|O\rangle$ appears to influence its CBC and BBC. We now show this is indeed the case, but before going on, we need to introduce the Fourier bra-ket entanglement (Fourier-BKE). Fourier-BKE is defined through coefficients $\{\lambda_{i}\}$, which are obtained from the Walsh-Hadamard transform [43] to the Schmidt coefficients $\{\gamma_{i}\}$ of BKE. $\{\gamma_{i}\}$ satisfies $\lambda_{i}\geq 0$ and $\sum_{i}\lambda_{i}^{2}=1$. Two sets of coefficients are related through $\Sigma=\sum_{i}\gamma_{i}|i\rangle\langle i|=2^{-n/2}\sum_{i}\lambda_{i}Q_{i}$ with $Q_{i}\in\{\pm 1\}\times\{I,Z\}^{\otimes n}$ (See SM B.2). The Rényi entropy on $\{\lambda_{i}\}$ is referred to as the Fourier-BKE Rényi entropy
@@
Detailed proofs can be found in SM D.2. Since CBC and BBC involve projecting $|O\rangle$ onto two different bases and BKE and Fourier-BKE are related by the Walsh-Hadamard transform, each of them also satisfies an entropic uncertainty principle [44].
@@
Detailed proofs can be found in SM B.2 and D.1.
Theorem 2 implies that BKE and Fourier-BKE set lower bounds for BBC and CBC. In this sense, CBC and BBC are like conjugate quantities that share the same importance in deciding SE. Also, due to the uncertainty principles in Theorem 3, CBC and BBC can neither be simultaneously small, nor can their lower bounds be simultaneously large. When the BKE of $O$ is low, $O$ must already have a lot BBC resource. Conversely, when the Fourier-BKE is low (BKE is high), $O$ must already have a lot CBC resource. Combined with Theorem 1, this suggests BKE can help diagnose which resource between CBC and BBC is more dominant for increasing SE of $O$. Since mixed states have a larger BKE compared to pure states, our results also indicate a noise-induced transition: as noise drives a pure state to a mixed state, the increase of the SE will change from governed by coherence to governed by magic.
We highlight that since Theorem 1 only sets lower bounds for CBC and BBC, BKE can lose diagnostic significance when CBC and BBC are already high in $|O\rangle$ since it no longer needs additional coherence resources for high SE. Nevertheless, we can view $U\otimes U^{*}|O\rangle=(U\otimes U^{*})(V\otimes V^{*})|O_{TEM}\rangle$ is prepared from a BKE-matching operator $O_{TEM}$. A BKE-matching operator has both CBC and BBC exactly reach the lower bounds in Theorem 2. BKE-matching operator also has the minimal uncertainties in Theorem 3 (See SM D.2). Examples of BKE-matching operators include tensor products of single-qubit operators picked from either $|i\rangle\langle j|$ or $P_{i}$. Due to the resulting equalities in Theorem 2, bra-ket bipartition precisely diagnoses the desired resources contained in $UV$ (See Lemma 7 in SM D.2).

### v3

# Bra-ket entanglement, an indicator bridging entanglement, magic, and coherence

@@ Abstract
Understanding the intricate interplay between distinct quantum resources is a fundamental prerequisite for rigorously characterizing the boundary between classical and quantum technologies. Among the vast landscape of quantum resources, entanglement, magic, and coherence have arguably attracted the most intense investigation. However, while universally recognized as the core drivers of quantum advantage, our understanding of their structural interplay remains fragmented and compartmentalized. In this work, we introduce an indicator called bra-ket entanglement (BKE) defined in the operator vectorization space to bridge all three quantum resources. Specifically, we show that BKE governs a resource dependence transition in the generation of entanglement: in the low-BKE regime, the growth of entanglement is dominated by coherence, largely independent of magic. However, as BKE increases, the dependence on coherence will gradually be replaced by a dependence on magic. Consequently, in the high-BKE regime, entanglement generation becomes dominated by magic, largely independent of coherence. These results are built on a series of new entropy-theoretic relations and are verified through numerical experiments. We also discuss implications of our results for the resource transitions in classical simulations of mixed states and marginal probabilities and for relating different classical simulation methods.
Introduction.—The ongoing quest to achieve and definitively prove quantum advantages [36, 37] relies fundamentally on our ability to understand and harness the underlying physical resources that remain inherently inaccessible to classical systems [14]. These unique quantum-mechanical resources serve a dual purpose in the landscape of modern quantum computing and information science [32]. On one hand, they act as the essential fuel that powers emerging quantum technologies [7, 3, 50, 29, 48, 47, 42]. On the other hand, these very same resources establish the fundamental limits for classically simulating quantum systems [34, 24, 1, 44, 12]. Identifying and quantifying these resources is crucial for determining which quantum systems are ”hard” to simulate and, consequently, where true quantum advantage lies.
Among the zoo of quantum features, three resources stand out as the pillars of quantum advantages: entanglement [28], magic [43], and coherence [41]. Entanglement, an intrinsic quantum correlation that plays a key role in quantum information applications such as quantum teleportation [7] and quantum communication [21]. In quantum computation, it also prevents efficient classical simulation by tensor network approaches [34, 16, 35] and has recently been shown to accelerate quantum simulations [50]. Magic, helps go beyond the stabilizer state, is a resource that enables universal quantum computation [10, 32] combined with Clifford circuits. Magic can be characterized as the level of superpositions of Pauli operators [30, 19], which set the bottleneck for efficient classical simulations via stabilizer-based simulation methods [24, 1] and Pauli-truncation based methods [39, 2, 4, 5]. Coherence, capturing the ability of a quantum system to exist in a coherent superposition of reference basis, is critical for quantum speedup due to the quantum interference effect in quantum algorithms [40] and enables Heisenberg-limit quantum metrology protocols [23]. Coherence also sets the complexity lower bound for classical simulation methods based on the Feynman path [12, 18].
Given the critical roles of these three resources, understanding their connections is a natural and necessary pursuit. Historically, however, they have been treated as distinct properties with different domains of application. For example, a stabilizer state [26] can possess maximal entanglement yet be efficiently simulatable due to zero magic [24, 1], while a product state with no entanglement can have both high magic [25] and high coherence [41]. The precise mechanism governing their interplay—and specifically, how one resource might be related to or constrain another—has remained unclear.
In this paper, we bridges these three resources by introducing an indicator which we call the bra-ket entanglement (BKE). Defined for general operators $O$ within the vectorized operator formalism, BKE is a conserved quantity under unitary evolution and can diagnose how the entanglement growth of the quantum evolution $UOU^{\dagger}$ is governed by the interplay between the coherence and magic. Here, $O$ can represent either a quantum state or a quantum observable, and $U$ represents the quantum evolution in the Schrödinger or in the Heisenberg picture, respectively.
Through BKE, we identify a resource dependence transition: when BKE of $O$ is low, entanglement growth is limited primarily by coherence (associated with Hadamard gates) and is independent of magic, explaining why stabilizer states can have high entanglement without magic. However, as BKE of $O$ increases, the dominant resource responsible for entanglement growth shifts from coherence to magic (associated with T gates), providing a deep underlying mechanism for the correlations between operator entanglement and Heisenberg magic observed in a recent study [20]. Therefore, we highlight the importance of this quantity of BKE in understanding and relating different quantum resources.
@@
Spatial vs bra-ket entanglement.—Throughout this work, we will adopt the vectorization picture to present our result, whose benifits will soon be clear. Let $O$ be an operator acting on the Hilbert space of $n$ qubits, hereafter denoted by ${\cal H}_{n}:=\mathbb{C}^{2^{n}}$, and let $|O\rangle:=\sum_{i,j}\,\langle i|O|j\rangle/\sqrt{\mathrm{Tr}[O^{\dagger}O]}\,|i\rangle|j\rangle$ be the corresponding normalized vector in the tensor product Hilbert space ${\cal H}_{n}\otimes{\cal H}_{n}$. Here, $O$ can be interpreted to represent either the density matrix or an observable.
@@
where $H_{\alpha}(\{p_{i}\}):=\log\left(\sum_{i}p_{i}^{\alpha}\right)/(1-\alpha)$ is the $\alpha$-order Rényi entropy [11]. We call $H_{SE,\alpha}\left(O\right)$ the space entanglement (SE) of the operator $O$. When $O$ is the density matrix of a pure quantum state, its SE is exactly twice the entanglement entropy of the state [28]. When the operator $O$ is the density matrix of a mixed state, the spatial entanglement is generally not a measure of mixed state entanglement (for example, it can be nonzero even for some separable states) [8, 27], but it is important because it determines the bond dimension, which in turn determines the complexity of classical simulations in the tensor network framework [34]. In the literature, $H_{SE,\alpha}\left(O\right)$ is sometimes known as the operator entanglement of operator $O$ [38].
@@
where $\{\gamma_{i}\}$ are the Schmidt coefficients in the decomposition $|O\rangle=\sum_{i}\gamma_{i}|O_{r,i}\rangle\otimes|O_{c,i}\rangle$, where $|O_{r,i}\rangle\}_{i}$ ($\{|O_{c,i}\rangle\}_{i}$) are orthogonal vectors acting on the Hilbert spaces associated to the row (column) indices (See SM B.2). For example, when $O$ is a rank-one operator like a pure state density matrix, $|O\rangle$ is a product state with zero BKE. In contrast, when $O$ is an $n$-qubit Pauli operator, $|O\rangle$ is proportional to a Bell state and has maximal BKE. In later sections, we will see how BKE relates the SE with two other quantum resources, namely magic and coherence. An important and nice feature of BKE is that it is a conserved quantity under unitary transformations (See SM B.2).
Magic and coherence in the vectorization picture.—Magic quantifies the deviation from stabilizer states, and the cost of simulating non-stabilizer states in the stabilizer framework [24, 1]. Here, we define a quantifier of magic in the vectorization picture. Our approach is to expand the vector $|O\rangle$ in the Bell basis with respect to the bipartition $(ab|cd)$ between bras and kets in the original operator $O$, and to define the Bell basis coherence (BBC) as the $\alpha$-order Rényi entropy
@@
Magic and coherence bound entanglement.—After the above definitions, we can now formally present our results. We fist show how the SE growth of $UOU^{\dagger}$ is bounded by magic and coherence. To do so, we define three free circuit families. The CBC-free (no coherence) circuit family $\mathcal{F}_{CBC}$ and BBC-free (no magic) circuit family $\mathcal{F}_{BBC}$ are defined as sets of unitaries $U$ satisfying $H_{CBC,\alpha}(O)=H_{CBC,\alpha}(UOU^{\dagger})$ and $H_{BBC,\alpha}(O)=H_{BBC,\alpha}(UOU^{\dagger})$, respectively, for arbitrary $|O\rangle$. Note that in the vectorization picture, we have $UOU^{\dagger}\rightarrow U\otimes U^{*}|O\rangle$. $\mathcal{F}_{CBC}$ is generated by diagonal phase and permutation gates on the computational basis, such as $\{T,S,CX\}$ or $\{CCX,S,CX\}$, while $\mathcal{F}_{BBC}$ consists of Clifford circuits generated by $\{H,S,CX\}$. The zero-coherence circuit family $\mathcal{F}_{zero}$ is the intersection of $\mathcal{F}_{CBC}$ and $\mathcal{F}_{BBC}$ with circuits constructed from $\{S,CX\}$ as examples. We discuss similarities and differences between $\mathcal{F}_{CBC}$ and $\mathcal{F}_{BBC}$ in SM E.1.
@@
Lemma 1.
For any set of orthogonal product basis $\{|\phi_{j}\rangle\}$ under the space bipartition in vectorization space, the decomposition $O=\sum_{j}c_{j}|\phi_{j}\rangle$ satisfies
$H_{\alpha}(\{|c_{j}|^{2}\})\geq H_{SE,\alpha}(O).$ (7)
The proof is given in SM D.1 with majorization relations [31].
In this work, we focus on two canonical choices: the computational basis $\{|i\rangle\!\langle j|\}$ and the Pauli basis $\{P_{j}/\sqrt{2^{n}}\}$, whose coefficient entropies correspond to CBC (coherence) and BBC (magic), respectively. Applying Lemma 1 to these two bases immediately yields the following theorem.
@@
Note that CBC and BBC are respectively sensitive to Hadamard and $T$ gates, which are responsible for coherence and magic generation, respectively. For a circuit $U$ with $N_{H}$ Hadamards and $N_{T}$ $T$ gates, we have (see SM C.1 and C.2):
@@
Therefore, in the absence of additional coherence and magic resources ($H$ and $T$), $G_{zero,\alpha}(O)$ is governed by the existing coherence and magic in $|O\rangle$. To generate higher entanglement, additional coherence and magic resources are needed, especially the smaller one. Theorem 1 explains why the stabilizer states can achieve maximal entanglement. Starting from $O=|0\rangle\langle 0|^{\otimes n}=\prod_{i=1}^{n}(I_{n}+Z_{i})/2$, $H_{CBC,\alpha}(O)=0$ severely limits $G_{zero,\alpha}(O)$, therefore, $H$ in $\mathcal{F}_{BBC}$ are necessary to increase RHS of Theorem 1. On the other hand, we have $H_{BBC,\alpha}(O)=n$ already reaches the largest achievable SE, meaning no need to introduce additional magic resources. Therefore, we have $0=G_{zero,\alpha}\left(|0\rangle^{\otimes n}|0\rangle^{\otimes n}\right)=G_{CBC,\alpha}\left(|0\rangle^{\otimes n}|0\rangle^{\otimes n}\right)$ while $G_{BBC,\alpha}\left(|0\rangle^{\otimes n}|0\rangle^{\otimes n}\right)=n$. Similar reasoning applies to $O=P$, with opposite behaviors.
BKE decides magic and coherence.—After showing that the entanglement growth is bounded by the smaller one between the coherence and the magic, now the second step is to understand what determines coherence and magic. From above, we observed that for $O=|0\rangle\langle 0|^{\otimes n}$ with minimal BKE, coherence is required to generate higher entanglement, while for $O=P$, with maximal BKE, magic is necessary to generate higher entanglement. Therefore, the BKE of $|O\rangle$ appears to influence its coherence and magic. We now show this is indeed the case, but before going on, we need to introduce the Fourier bra-ket entanglement (Fourier-BKE). Fourier-BKE is defined through coefficients $\{\lambda_{i}\}$, which are obtained from the Walsh-Hadamard transform [33] to the Schmidt coefficients $\{\gamma_{i}\}$ of BKE. $\{\gamma_{i}\}$ satisfies $\lambda_{i}\geq 0$ and $\sum_{i}\lambda_{i}^{2}=1$. Two sets of coefficients are related through $\Sigma=\sum_{i}\gamma_{i}|i\rangle\langle i|=2^{-n/2}\sum_{i}\lambda_{i}Q_{i}$ with $Q_{i}\in\{\pm 1\}\times\{I,Z\}^{\otimes n}$ (See SM B.2). The Rényi entropy on $\{\lambda_{i}\}$ is referred to as the Fourier-BKE Rényi entropy
@@
Detailed proofs can be found in SM D.2. Theorem 2 implies that BKE and Fourier-BKE set lower bounds for BBC and CBC. In this sense, coherence and magic are like conjugate quantities that share the same position in deciding entanglement growth. Since CBC and BBC involve projecting $|O\rangle$ onto two different bases and BKE and Fourier-BKE are related by the Walsh-Hadamard transform, each of them also satisfies an entropic uncertainty principle [17].
@@
Detailed proofs can be found in SM B.2 and D.1. Due to the uncertainty principles in Theorem 3, CBC and BBC can neither be simultaneously small, nor can their lower bounds be simultaneously large.
Finishing the picture.—Combining Theorems 13, the relation between entanglement, magic, and coherence can now be summarized in Fig. 2. When BKE is low, Theorem 2 forces BBC to be large while CBC can be small, so CBC becomes the bottleneck for entanglement growth and coherence ($H$ gates) is the dominant resource for increasing SE. Conversely, when BKE is high (Fourier-BKE low), CBC is forced large while BBC can be small, making magic ($T$ gates) the dominant resource.
This picture becomes fully quantitative for operators that saturate the inequalities in Theorems 2 and 3, which we call BKE-matching operators (See formal definition in SM D.2). For such operators with $H_{\mathrm{BKE},1/2}(O)=r$, the CBC and BBC values are exactly determined: $H_{\mathrm{CBC},\alpha}(O)=r$ and $H_{\mathrm{BBC},\alpha}(O)=n-r$ (see Lemma 8 in SM D.2). Combined with the gate-counting bounds in Eqs. (11)–(12), this yields a precise gate-complexity characterization.
Figure 2: Relating (space) entanglement to CBC (coherence) and BBC (magic) via BKE. When BKE of $O$ is low, the SE is bounded by CBC, and therefore increasing SE requires introducing coherence resources. When BKE of $O$ is high, the SE is bounded by BBC, and therefore increasing SE requires introducing magic resources. When BKE is medium, both magic and coherence can relate SE, but not as strongly as in the extreme points.
Corollary 1.
For a BKE-matching operator $O$ with fixed bra-ket entanglement $H_{\mathrm{BKE},1/2}(O)=r$, if a circuit $U$ composed of $\{H,S,\mathrm{CNOT},T\}$ gates with $N_{T}$ number of $T$-gate and $N_{H}$ number of $H$-gate satisfies $H_{\mathrm{SE},\alpha}(U^{\dagger}O\,U)\geq\eta$, then $N_{T}\geq\eta-(n-r),N_{H}\geq\eta-r$.
The proof follows from Lemma 8 in SM D.2. When $r=0$ (e.g. $O=|0\rangle\!\langle 0|^{\otimes n}$), the $N_{T}$ bound becomes trivial while $N_{H}\geq\eta$ dominates, recovering the coherence-limited regime discussed after Theorem 1: stabilizer states achieve maximal entanglement without magic. When $r=n$ (e.g. $O=P$), the situation reverses and $N_{T}\geq\eta$ dominates, explaining the entanglement–magic correlation observed in Heisenberg evolution of Pauli operators [20]. At intermediate BKE, both bounds impose comparable constraints, so neither coherence nor magic alone capture the tensor-network complexity.